I am trying to convert these words into an equation:
Find three consecutive odd integers such that the sum of three times the first and twice the second is 26 more than three times the third ?
Can you please help?
Thank you.
: I am trying to convert these words into an equation:
Find three consecutive odd integers such that the sum of three times the first and twice the second is 26 more than three times the third ?
Can you please help?
Thank you.

Answer by stanbon(18754)

Find three consecutive odd integers such that the sum of three times the first and twice the second is 26 more than three times the third ?
------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
-------------
3(2x-1)+2(2x+1) = 3(2x+3)+26
6x-3 + 4x+2 = 6x+9+26
10x -1 = 6x + 35
4x = 36
x = 9
1st: 2x-1 = 17
2nd: 19
3rd: 21
===================
Cheers,
Stan H.